OpenStudy (anonymous):

solve for radical equation. √5x-4=6

OpenStudy (anonymous):

√5x-4=6 √5x=6+4 √5x=10 \[x = \frac{10}{\sqrt 5} =\frac{5 \times 2}{\sqrt 5}=\frac{\sqrt 5 \times \sqrt 5 \times 2}{\sqrt 5}\] \[\rightarrow x =\frac{1 \times \sqrt 5 \times 2}{1}= 2\sqrt 5\] Thus \[x=2\sqrt5\] is the solution of this equation

OpenStudy (anonymous):

@DatDude